OpenStudy (anonymous):

What is the value of b in the triangle shown below? –4 in. 4 in. ±4 in. There is no solution.

OpenStudy (anonymous):

here is the photo

OpenStudy (saifoo.khan):

Use the area formula, Area= 0.5 * b *h 24 = 0.5* b*3b Now, solve for b.